OpenStudy (fernandez555):

medal+fan. Simplify i^22 A. 1 B. −1 C. −i D. i

OpenStudy (usukidoll):

well we could split this up into small i^2 s because i^2 is equal to -1

OpenStudy (usukidoll):

I know I'm doing this the long way... (i^2)(i^2)(i^2)(i^2)(i^2)(i^2)(i^2)(i^2)(i^2)(i^2)(i^2) (-1)(-1)(-1)(-1)(-1)(-1)(-1)(-1)(-1)(-1)(-1) and then we find some pairs to make it faster.. I got 5 pairs of (-1)s

OpenStudy (usukidoll):

(-1)(-1)(-1)(-1)(-1)(-1)(-1)(-1)(-1)(-1)(-1) (1)(1)(1)(1)(-1) = -1

OpenStudy (usukidoll):

ok it is -1 and I figured out a shortcut. we know that i^2 = -1 so we could modify the problem a bit (i^2)^11 because 11 x 2 is 22 via distributive exponential rules since i^2 = -1 (-1)^11 and we have an odd exponential so the result will be negative.... -1

OpenStudy (usukidoll):

B

OpenStudy (anonymous):

i^1 = i i^2= -1 i^3= -i i^4= 1 i^5 = i It repeats every 4